OpenStudy (jmartinez638):

What are the values of k for which:

OpenStudy (jmartinez638):

\[\int\limits_{-3}^{k}x^2dx=0\]

OpenStudy (anonymous):

one value should be more or less obvious do you know it ?

OpenStudy (jmartinez638):

3

OpenStudy (jmartinez638):

No, actually. not three

OpenStudy (anonymous):

hmm no i don't think 3 works this function is always postie so any integral will also be positive unlesss

OpenStudy (anonymous):

"positive"

jimthompson5910 (jim_thompson5910):

k = 3 would work only if the function was odd but instead, you use this rule \[\Large \int_{k}^{k}f(x)dx = 0\]

OpenStudy (jmartinez638):

It'd have to be -3

jimthompson5910 (jim_thompson5910):

the integral I wrote represents the area under the curve from x = k to x = k, which is an infinitely thin rectangle with 0 area

jimthompson5910 (jim_thompson5910):

yes, k = -3

OpenStudy (jmartinez638):

And that would have to be the only value.

jimthompson5910 (jim_thompson5910):

yes because x^2 is always positive (well the nonzero x values anyway) and never dips below the x axis

OpenStudy (jmartinez638):

Makes sense.

jimthompson5910 (jim_thompson5910):

Here is another way to solve Let \[\Large g(x) = \int x^2dx\] You'll find, after integrating, that \[\Large g(x) = \frac{x^3}{3}+C\] where C is any constant ------------------------- \[\Large g(x) = \frac{x^3}{3}+C\] \[\Large g(k) = \frac{k^3}{3}+C\] ------------------------- \[\Large g(x) = \frac{x^3}{3}+C\] \[\Large g(-3) = \frac{(-3)^3}{3}+C\] \[\Large g(-3) = \frac{-27}{3}+C\] ------------------------- so, \[\Large \int_{-3}^{k}x^2dx = g(k) - g(-3)\] \[\Large \int_{-3}^{k}x^2dx = \left(\frac{k^3}{3}+C\right)- \left(\frac{-27}{3}+C\right) = 0\] I'm going to skip a bunch of steps, but solving `(1/3)k^3+C - (-27/3)-C = 0` for k leads to k = -3. Technically there are 2 other solutions, but they are not real numbers. So we can ignore them.